OpenStudy (anonymous):

if sin20=x express Tan 20 in terms of x

Parth (parthkohli):

Hint:\[{\sin\theta\over \cos \theta} = \tan \theta\]

OpenStudy (anonymous):

how do u find cos20

Parth (parthkohli):

Using a calculator?

OpenStudy (anonymous):

i need 2 express tan 20 in terms of x im stuck on how to express cos20 in terms of x

Parth (parthkohli):

\[\cos \theta = \sqrt{1 - \sin^2 \theta}\]

OpenStudy (anonymous):

thanx so the final answer is....................

OpenStudy (anonymous):

the final answer is tan20=x/(sqrt(1-x^2))